OpenStudy (anonymous):

let _g_ be defined as _g_=2g^2-2 for all values of g. If _x_=x^2, what is the value of x? @whpalmer4

OpenStudy (whpalmer4):

Isn't that saying \[x^2 = 2x^2-2\]

OpenStudy (anonymous):

is it \[\sqrt{2}\]?

OpenStudy (whpalmer4):

Let's find out! \[G(\sqrt{2}) = 2(\sqrt{2})^2-2 = 2(\sqrt{2}*\sqrt{2}) - 2 = 2*2-2 = 4-2=2\]That's what we wanted.